### v1

# AstaGen: Empirical Evaluation of Automated SATD Taxonomy Generation with LLMs

Abstract
Technical debt refers to suboptimal code that degrades software quality. When developers intentionally introduce such debt, it is called self-admitted technical debt (SATD). Since SATD hinders maintenance, identifying its categories is key to uncovering quality issues. Traditionally, constructing such taxonomies requires manually inspecting SATD comments and surrounding code, which is time-consuming, labor-intensive, and often inconsistent due to annotator subjectivity.
This study presents AstaGen, an initial step toward automating SATD taxonomy generation using large language models (LLMs). Given a comment and its surrounding code, AstaGen first generates a concise explanation for each SATD comment, then incrementally generates and updates categories to construct a taxonomy. We evaluate AstaGen on SATD datasets from three domains: quantum software, smart contracts, and machine learning. It successfully recovers domain-specific categories reported in prior work, such as Layer Configuration in machine learning. Compared to a naive use of an LLM, AstaGen produces more consistent category assignments due to its explanation-driven, iterative design. It also completes taxonomy generation in under two hours and for less than one USD, even on the largest dataset. These results suggest that while full automation remains challenging, AstaGen is able to support semi-automated taxonomy construction. Furthermore, our work opens up avenues for future work, such as automatic taxonomy generation in other areas.
Index Terms:
Self-Admitted Technical Debt, Large Language Model, Automated Taxonomy Generation
I Introduction
Developers sometimes introduce suboptimal code that degrades software quality; this is known as technical debt [1]. In particular, when developers intentionally insert such code due to constraints such as budget or schedule, it is referred to as Self-Admitted Technical Debt (SATD) [2]. In practice, developers explicitly mark SATD via code comments.
SATD can negatively affect software maintenance and operation. Consequently, identifying SATD categories and addressing them can improve code quality [3, 4]. Previous studies therefore developed reusable SATD taxonomies that systematically organize these debts [3, 5, 6]. For example, Maldonado and Shihab [3] defined a taxonomy for Java projects with five categories: design, requirement, defect, test, and documentation. Ishimoto et al. [5] extended this idea to quantum software by incorporating quantum-specific factors, and Ebrahimi et al. [6] developed a taxonomy for smart contracts. Similarly, Obrien et al. [7] proposed a taxonomy tailored to machine learning systems.
While SATD taxonomies help in detecting and remedying debt, creating them is costly because it requires at least two researchers to manually analyze SATD comments [3, 5, 6, 7]. While existing taxonomies can be reused within the same domain (e.g., Java), new domains (e.g., machine learning software) require building a new taxonomy. Moreover, manual analysis introduces the subjectivity of researchers and suffers from poor reproducibility; different teams may produce different taxonomies. Therefore, automating SATD taxonomy generation remains a critical research challenge.
To this end, we propose a large language model (LLM)-driven approach for SATD taxonomy generation, which we call AstaGen (Automated SATD TAxonomy GENeration). Traditional clustering techniques and learning-based classifiers struggle to (a) understand the issues raised by SATD comments and (b) generate new categories for them, even though they can group similar comments or fit them into existing categories. To address these limitations, we design AstaGen as a two-phase LLM-based approach that (a’) generates a concise explanation of the issue raised by each SATD comment, and (b’) iteratively generates and updates categories based on these explanations to construct a coherent taxonomy.
We evaluated AstaGen across three domains (i.e., quantum software [5], smart contracts [6], and machine learning software [7]) using human-defined taxonomies as references. AstaGen successfully generated domain-specific categories that were semantically similar to those defined by humans, such as Layer Configuration in machine learning. Compared to a naive use of an LLM, AstaGen generated more consistent taxonomy due to its explanation-driven, iterative design. In addition, AstaGen completed taxonomy generation in under two hours and at a cost of less than one USD, even for the largest dataset containing 448 comments, highlighting its efficiency compared to manual approaches.
Our main contributions are as follows: (1) We propose AstaGen, an automated SATD taxonomy generation approach using LLMs. (2) AstaGen can generate domain-specific categories that are semantically similar to those defined by humans. (3) AstaGen generates taxonomies that are more aligned with human-defined ones compared to a naive use of an LLM. (4) AstaGen is a time- and cost-friendly approach; it runs in under two hours and for less than one USD. (5) We suggest practical use cases for AstaGen in semi-automated taxonomy generation supported by our empirical results.
II Related Work
II-A Construction of SATD Taxonomies
To support the understanding of SATD, prior studies have investigated both the types and frequencies of SATD comments in source code [3, 4, 5, 6, 8]. In particular, constructing a SATD taxonomy (i.e., classification scheme for SATD comments) has been recognized as essential for achieving a deeper understanding of technical debt in software projects.
Constructing such a taxonomy involves two main tasks. First, researchers identify SATD-related comments by applying keyword-based filters, typically regular expressions for markers such as todo and fixme [2, 3], followed by manual validation to confirm their relevance. Second, they categorize the filtered comments by manual inspection: at least two researchers read the comments, propose category labels, and reconcile their labels through discussion until consensus is reached. For example, Ishimoto et al. [5] engaged three researchers to manually inspect 202 comments, identifying 193 SATD instances (including 88 quantum-specific cases) to construct a domain-tailored taxonomy. Similarly, Ebrahimi et al. [6] had two authors manually review 287 SATD comments to extend existing taxonomies to smart contracts.
Although manual taxonomy construction is crucial for understanding SATD, it suffers from three major limitations. (1) It is time-consuming: Xiao et al. [8] required 120 person-hours to inspect 500 SATD comments. (2) It does not scale: Kashiwa et al. [9] analyzed only 375 of 15,671 SATD comments to curb manual effort. (3) It depends heavily on the prior knowledge and experience of the annotators: when independently labeling 50 comments, Ebrahimi et al. [6] observed 12 inconsistencies and relied on the second author, who had greater smart contract expertise, to resolve them. These issues stem from the manual nature of taxonomy construction. Therefore, we propose an initial approach toward automating SATD taxonomy generation using LLMs.
Figure 1: SATD detection, classification, and taxonomy generation.
II-B Automated SATD Detection and Classification
Prior studies have automated both the detection of SATD in source code comments [10, 11, 12] and the classification of detected SATD based on existing taxonomies [13, 14]. Figure 1 illustrates these two tasks. Chen et al. [13] used XGBoost to automatically classify three SATD types: defect, design, and implementation. Sheikhaei et al. [14] applied both fine-tuning and in-context learning to LLMs to detect and classify SATD comments into predefined categories, achieving higher accuracy than traditional machine learning baselines. These prior studies overlooked the task of SATD taxonomy generation (Figure 1). Our work aims to fill this gap by automating SATD taxonomy generation, a novel task that involves understanding the intent of each SATD to generate new categories.
II-C LLMs as a Replacement for Manual Annotation
Some prior studies have explored the potential of LLMs to replace human annotators in software engineering tasks [15, 16, 17]. Ahmed et al. [15] examined whether LLMs can replace human annotators in software engineering tasks. Their study involves ten annotation tasks, including assessing the quality of code summaries, all of which are framed as multiple choice questions with predefined answer options. Their results indicate that LLMs can replace human annotators for some tasks with comparable agreement levels. Our target task, SATD taxonomy generation, is disjoint from theirs, as it requires generating category labels rather than selecting from predefined ones. Indeed, they left such tasks (i.e., open coding) to future work. We push this direction forward through the task of SATD taxonomy generation.
Figure 2: Workflow of AstaGen. The robot icons above the arrows indicate the steps in which LLMs are utilized.
III AstaGen: Automated SATD Taxonomy Generation by LLMs
III-A Overview
Figure 2 illustrates the overall workflow of AstaGen.
Input and Output. The input consists of a set of SATD comments and their corresponding source code. The output is a SATD taxonomy generated by LLMs based on these inputs. The resulting classification scheme is structured as a two-level tree. We refer to the upper and lower levels of this hierarchy as the main categories and subcategories, respectively. This structure aligns with most existing studies on SATD taxonomy construction [4, 6, 9, 5], which adopt trees with at least two hierarchical levels. In addition to generating the taxonomy, AstaGen also assigns each input comment to a specific main category and subcategory in the resulting taxonomy.
Workflow. AstaGen consists of two phases: (1) generating explanations for SATD, and (2) generating and updating categories at each level of the taxonomy. The two-phase design aims to efficiently generate the SATD taxonomy while accommodating the context length limitations of LLMs. In the first phase, concise explanations serve as compressed representations of SATD comments and their corresponding source code. In the second phase, the taxonomy is iteratively updated after each batch of comments, helping LLMs maintain consistency for generated categories. This iterative refinement mimics the common practice in human-driven taxonomy generation, where categories are periodically reviewed after processing a certain number of comments [4, 5, 8]. The following subsections describe each phase in detail.
III-B Phase 1: Generating Explanations for SATD
In this phase, the LLM generates a concise explanation (typically 2–3 sentences) for each SATD comment, describing its content and context. Prior studies have shown that the surrounding source code provides useful context for understanding SATD [14, 12]. Therefore, our prompts include not only the comment itself but also the relevant source code. Specifically, we include the entire file in which the comment appears, enabling the LLM to generate explanations that reflect both the comment and the associated code issue.
III-C Phase 2: Generating/Updating Categories at Each Level
In this phase, the LLM constructs a SATD taxonomy consisting of main and subcategories. The taxonomy is built hierarchically: main categories are generated first, followed by subcategories. Each level involves two steps—category generation and update—applied iteratively in batches, each containing a fixed number of SATD explanations. This top-down structure allows the LLM to focus on a single level of abstraction at a time. As batching proceeds, the list of categories is continuously updated until all comments are processed. The output of this phase is a complete two-level taxonomy derived after processing all batches at the subcategory level. The following describes the generation and update steps.
III-C1 Generation Step
In this step, the LLM generates a category name of approximately 1–3 words for each comment included in the batch, based on its corresponding SATD explanation. All explanations in the current batch are enumerated and presented together in a single prompt. The LLM then produces a category name for each explanation individually.
III-C2 Update Step
In this step, the categories generated in the current batch are integrated with those from previous batches to update the category list.
First, the LLM integrates newly generated categories with those from previous batches. The prompt includes category names and their associated explanations: all explanations from the current batch, and up to three randomly sampled explanations per category from previous batches. The LLM is instructed to merge semantically similar categories based on this information. By comparing category names before and after integration, we obtain a mapping of name changes.
Second, the mapping is used to update the category list. This ensures consistency across all processed comments. Since this update involves only applying the name-change mapping, it does not require further use of the LLM.
Finally, AstaGen outputs a category assignment for each comment. These categories are based on the SATD taxonomy generated by AstaGen, which is represented as a two-level taxonomy tree consisting of main categories and subcategories.
IV Study Design
IV-A Research Questions
  1. RQ1
    (Category name alignment) Can AstaGen generate category names that are semantically similar to those defined by humans?
  2. RQ2
    (Taxonomy content alignment) Does AstaGen achieve better alignment with human-defined taxonomies than a naive use of an LLM, in terms of comment assignment and category granularity?
  3. RQ3
    (Efficiency) How much time and cost does AstaGen require?
RQ1 and RQ2 aim to compare the taxonomy generated by AstaGen with that created by humans. RQ1 focuses on the similarity of category names, while RQ2 examines the consistency of the comments assigned to each category. RQ3 assesses the practicality of AstaGen by measuring the execution time and monetary cost (i.e., API fees for LLMs) required to generate the SATD categories.
IV-B Studied Datasets
We use SATD comments and corresponding taxonomies from prior studies as our datasets. The SATD comments are input to AstaGen, while the human-defined taxonomies serve as references for evaluation.
We collected papers that constructed SATD taxonomies from two major digital libraries: IEEE Xplore and the ACM Digital Library. Specifically, we used the following query to search for papers whose abstracts contained specific keywords: ("SATD" OR "self-admitted technical debt") AND ("taxonomy" OR "category" OR "coding"). The part before AND is intended to capture papers related to SATD, while the part after AND aims to capture those involving taxonomy or category construction. After removing duplicates, this query returned 79 papers.111 Search conducted on January 8, 2025; reconfirmed unchanged on May 16, 2025.
From these, we selected papers that met both of the following criteria: (1) the taxonomy was manually constructed, and (2) the corresponding data was available. Two authors independently screened the papers, resulting in five candidate papers [4, 18, 7, 6, 5]. The first author then thoroughly reviewed the data of these five studies. Papers were excluded if the data were incomplete (e.g., only top-level categories were available) or inconsistent with the taxonomy in the paper (e.g., category names differed from those reported in the paper). As a result, the following three datasets were selected for use in this study:
  • Quantum Software (QS) [5]: 88 SATD comments from Qiskit[19]-dependent projects; includes 4 main and 9 subcategories.
  • Smart Contracts (SC) [6]: 190 SATD comments from Solidity-based contracts; includes 6 main and 26 subcategories.
  • Machine Learning Software (ML) [7]: 448 SATD comments from ML-related GitHub repositories; includes 9 main and 23 subcategories.
IV-C Selected Models
In all experiments, we used Deepseek-V3 [20] as the LLM, with the default temperature setting of 1.0. Since the output of the LLM can vary across runs, we repeated all experiments ten times to assess the stability of the generated results.
In RQ1, we evaluate the semantic similarity between pairs of category names—those defined by humans and those generated by AstaGen. We encode each category name into a 384-dimensional vector using the pre-trained sentence embedding model all-MiniLM-L6-v2, [^2] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and compute the cosine similarity between the resulting vectors. We selected a sentence-level embedding model because category names often consist of multiple words, and such models are better suited for capturing their overall semantics.
IV-D Prompt Design
In Phase 1 of AstaGen (Section III-B), we limited the source code context given to the LLM to 2,000 lines to comply with its context length limit. If the file exceeded this limit, we included only the 1,000 lines before and after the SATD comment. We used the line numbers available in the original datasets to accurately locate each comment.
In Phase 2 of AstaGen (Section III-C), we used a batch size of 20. Specifically, SATD explanations were grouped into batches of 20, and each batch was processed with a separate prompt.
IV-E Evaluation Metrics
This section describes the evaluation metrics used in each RQ. For each dataset and taxonomy level, let $\mathcal{H}$ and $\mathcal{M}$ be the sets of human-defined and AstaGen-generated category names, respectively. We denote individual category names from these sets as $H_{i}\in\mathcal{H}$ and $M_{j}\in\mathcal{M}$$\mathcal{M}$ stands for “machine-generated,” in contrast to $\mathcal{H}$ for “human-defined.”
IV-E1 RQ1
We compute the cosine similarity between the sentence embeddings for $H_{i}$ and $M_{i}$, denoted as $sim(H_{i},M_{j})$. We then compute $top\_sim(H_{i})=\max_{j}sim(H_{i},M_{j})$, which represents the highest similarity between $H_{i}$ and any generated name. This score quantifies how closely AstaGen can generate a category name for each human-defined category. We visualize the distribution of $top\_sim(H_{i})$ for each dataset and level to assess the alignment between human-defined and AstaGen-generated category names.
Figure 3: Example Sankey diagram of human-defined and AstaGen-generated categories (Dataset: QS, Level: Main, Run ID: 0).
IV-E2 RQ2
We evaluate whether similar comments are grouped into similar categories by humans and AstaGen. Unlike classification problems, $H_{i}$ and $M_{j}$ are drawn from different sets (i.e., $\mathcal{H}$ and $\mathcal{M}$), making conventional precision and recall inapplicable. To address this, we define a best-match precision and recall for each human-defined category. Let $C_{ij}$ denote the number of comments assigned to both $H_{i}$ (by humans) and $M_{j}$ (by AstaGen). The matrix $C$ enables the construction of a Sankey diagram (e.g., Figure 3) that visualizes the flow of comments between human-defined and AstaGen-generated categories. [^4] The thickness of each flow from $H_{i}$ to $M_{j}$ is proportional to $C_{ij}$. For each $H_{i}$, we define its best-matched category as $M_{j^{*}}$, where $j^{*}=\operatorname{argmax}_{j}C_{ij}$. Using this best match, we compute:
$\text{precision}(H_{i})=\frac{C_{ij^{*}}}{\sum_{i=1}^{|\mathcal{H}|}C_{ij^{*}}},\quad\text{recall}(H_{i})=\frac{C_{ij^{*}}}{\sum_{j=1}^{|\mathcal{M}|}C_{ij}}.$
Here, precision reflects how exclusively the best-matched category corresponds to the human-defined category $H_{i}$, while recall indicates how well the comments in $H_{i}$ are covered by the best-matched category. We also report the F1 score as the harmonic mean of precision and recall.
Our proposed metrics quantify the correspondence between human-defined and AstaGen-generated categories as visualized in the Sankey diagram. For example, in Figure 3, the human-defined category Backend is best matched with Quantum System Integration from AstaGen, yielding a precision of 0.45 and a recall of 0.76. As shown in Figure 3, Backend category is largely covered by Quantum System Integration, indicating high recall. On the other hand, since Quantum System Integration also includes comments from other human-defined categories, its precision is low.
While these metrics quantify the degree of comment-level alignment between human-defined and AstaGen-generated categories, they are sensitive to the total number of categories. For instance, if AstaGen generates only a few categories, recall can become high because all the comments in a human-defined category may fall under a single generated category. Therefore, we also report the difference in the number of generated and human-defined categories: $|\mathcal{M}|-|\mathcal{H}|$. A smaller absolute value indicates that the number of categories generated by AstaGen more closely matches the human-defined categories. We refer to this aspect as category granularity.
IV-E3 RQ3
For each dataset, we calculate the average execution time and cost of running AstaGen ten times. In AstaGen, the cost only arises from the use of the DeepSeek-V3 API. During each run, our implementation records the total number of input and output tokens, and estimates the API cost based on the official pricing table of DeepSeek-V3.555https://api-docs.deepseek.com/quick_start/pricing, as of May 20, 2025
V Results
Figure 4: Distribution of $top\_sim(H_{i})$ scores across datasets and taxonomy levels. The annotated examples (e.g., Algorithm, Layer Configuration) correspond to the top-5 highest similarity scores across all data points.
V-A RQ1: Category Name Alignment
Figure 4 shows the distribution of $top\_sim(H_{i})$ scores. The x-axis denotes the dataset, and colors indicate the taxonomy level. The mean $top\_sim$ scores for the QS, SC, and ML datasets are 0.36, 0.32, and 0.42, respectively.
AstaGen can generate domain-specific category names that are similar to human-defined ones. As highlighted in Figure 4, all points with scores above $\geq 0.8$ correspond to domain-specific categories (e.g., Layer Configuration in ML).
High-similarity categories are more likely to be domain-specific, whereas low-similarity categories tend to be generic. To examine the relationship between $top\_sim(H_{i})$ and the type of category, two authors independently labeled each human-defined category in the top and bottom 10% of $top\_sim(H_{i})$ as either domain-specific or generic, counting only those with agreement. Among the top 10% (77 categories), 46 ($\simeq$ 60%) were domain-specific, supporting the earlier observation. In the bottom 10%, 53 ($\simeq$ 69%) were labeled as generic. These results suggest that AstaGen more effective at generating consistent names for domain-specific categories, while generic categories are harder to match semantically.
Answer to RQ1: AstaGen tends to generate domain-specific category names that are semantically similar to human-defined names (e.g., Layer Configuration in ML). In contrast, low-similarity cases were predominantly generic (69% of 77 categories).
TABLE I: Comparison of AstaGen and its naive variant.
Data. Lv. Pre. Rec. F1 $|\mathcal{M}|-|\mathcal{H}|$
A N A N A N A N
QS Main 0.43 0.25 0.59 0.77 0.50 0.37 0.3 0.5
Sub 0.33 0.16 0.39 0.68 0.36 0.26 1.9 1.5
SC Main 0.20 0.27 0.63 0.80 0.31 0.41 -2.0 -1.6
Sub 0.19 0.14 0.56 0.79 0.29 0.25 -8.2 -15.5
ML Main 0.15 0.12 0.45 0.87 0.22 0.20 1.6 -6.4
Sub 0.17 0.11 0.36 0.67 0.23 0.19 6.1 -13.5
A: AstaGen, N: Naive. Bold indicates the better value.
V-B RQ2: Taxonomy Content Alignment
Table I presents the mean precision, recall, F1 scores, and differences in the number of categories for each dataset and taxonomy level. Although these metrics capture two aspects (i.e., comment assignment accuracy and category granularity), there are no guidelines for interpreting their absolute values. Therefore, we introduce a baseline, the naive variant, which generates the taxonomy using an LLM without Phases 1 and 2 of AstaGen. This baseline allows us to evaluate the extent to which AstaGen improves alignment with the human-defined taxonomy compared to a naive use of LLMs.
AstaGen achieves higher F1 scores than the naive variant in five out of six cases, indicating a better balance between precision and recall. This result suggests that core components of AstaGen, which the naive variant lacks, are effective in improving the alignment of comment assignments between human-defined and AstaGen-generated categories.
AstaGen also better matches the granularity of human-defined taxonomies in four out of six cases. In four out of six cases, AstaGen exhibits a smaller absolute difference than the naive variant, suggesting that its taxonomy size more closely matches human taxonomies. In contrast, the naive variant often produces fewer categories, which likely contributes to its higher recall. These findings indicate that AstaGen outperforms the naive variant by producing taxonomies that are not only more accurate in assignment but also more comparable in structural granularity. Hence, AstaGen generates similar categories to human-defined ones compared to the naive variant.
Answer to RQ2: AstaGen achieves higher F1 scores than the naive variant in five out of six cases. It also generates a number of categories more similar to human-defined taxonomies in four out of six cases. Hence, AstaGen generates categories that are more similar to human-defined ones compared to the naive variant. However, the absolute values of precision and recall remain between 0.152 and 0.638, indicating that there is still room for improvement in comment-level alignment.
TABLE II: Average execution time and cost of AstaGen for each dataset.
Dataset (Size) Avg. Time [min] Avg. Cost [USD]
QS (88) 21.60 0.142
SC (190) 52.98 0.614
ML (448) 109.50 0.813
V-C RQ3: Efficiency (Time and Cost)
Table II shows the average execution time and cost of AstaGen for each dataset, averaged over ten runs.
Compared to prior manual efforts, AstaGen achieves significantly reduces both time and cost. For example, Ebrahimi et al. [6] reported spending 57 person-hours just to construct the initial taxonomy, while Xiao et al. [8] required 120 person-hours to manually inspect 500 SATD comments. In contrast, AstaGen categorizes hundreds of comments in under two hours at a cost of less than one USD. These substantial efficiency gains highlight the practicality of using LLMs for large scale SATD taxonomy generation.
Answer to RQ3: AstaGen is efficient in terms of both time and cost. Even for the largest dataset we used (i.e., the ML dataset with 448 comments), it took only 109.50 minutes and cost just 0.813 USD on average.
VI Use Case: Semi-Automated Taxonomy Generation
Although AstaGen may not fully automate SATD taxonomy generation—as RQ1 and RQ2 showed that its alignment with human-defined taxonomies was not always high—it remains a practical assistant for reducing human effort. We propose two use cases for AstaGen: (1) as a virtual annotator that can replace some human annotators in collaborative taxonomy construction, and (2) as an initial step that provides category candidates to support annotators. These use cases support semi-automated taxonomy generation. By incorporating these human-in-the-loop workflows, it becomes possible to balance the efficiency of automation (as demonstrated in RQ3) with the quality of the resulting taxonomy through human review.
Notably, AstaGen has demonstrated the ability to generate domain-specific categories. This makes it especially valuable when researchers must construct a taxonomy in an unfamiliar domain. By incorporating AstaGen as one of the contributors, researchers can accelerate the early stages of taxonomy construction. In addition to improving AstaGen, evaluating these use cases remains future work.
VII Threats to Validity
Construct Validity. In RQ2, we evaluate the alignment of AstaGen-generated taxonomies with human-defined ones using best-match precision and recall. This evaluation is conducted in a relative manner using the naive variant, and therefore lacks an absolute evaluation. Further investigation is needed to examine the correlation between these metric values and human preferences for each generated category. This limitation is part of a broader challenge known as alignment evaluation for LLMs [21]. Finally, we treat the final human-defined categories from the original studies as references, although these may be inconsistent or incomplete.
Internal Validity. The outputs of AstaGen depend on the behavior of the underlying LLM (DeepSeek-V3), which is stochastic by nature. To mitigate variability, we repeated each run ten times. Still, some outcomes may be sensitive to prompt design or batch composition, particularly in Phase 2, where category merging occurs.
External Validity. Our evaluation used SATD datasets from three domains—quantum software, smart contracts, and machine learning software. While these cover diverse domains, they may not generalize to all types of software projects.
VIII Conclusion
We present AstaGen, an initial step toward automating SATD taxonomy generation using LLMs. Across three datasets, AstaGen (i) recovered domain-specific categories reported in prior studies (e.g., Layer Configuration in the ML dataset), (ii) outperformed a naive baseline in comment assignment by up to 0.13 F1, and (iii) completed the entire process for 448 comments in under two hours at a cost of less than one USD even for the largest dataset.
Our findings suggest that fully automating SATD taxonomy generation remains a challenging task. Nevertheless, employing AstaGen as a virtual annotator or as an initial step to propose candidate categories may help reduce human effort. In future work, we plan to explore semi-automated, human-in-the-loop workflows and examine the trade-offs between the degree of automation and taxonomy quality. Furthermore, extending automated taxonomy generation beyond SATD to other open coding tasks (e.g., a taxonomy of root causes of fuzzing build failure [22]) remains an promising direction for future research.
Acknowledgment
We gratefully acknowledge the financial support of: (1) JSPS for the KAKENHI grants (JP24K02921, JP25K03100); (2) Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant NumberJPMJAP2415, and (3) the Inamori Research Institute for Science for supporting Yasutaka Kamei via the InaRIS Fellowship.